# Semigroups uniquely determined by one-sided identity and zero sets

### 31 Mar 2026 v3

@@ Abstract
classification: 20M10
keywords: identity set; zero set; idempotent; stabilized semigroup; commutative-rectangular band
@@ 3 Identity and Zero Relations and Sets
Let $S$ be a semigroup. For every idempotent $e\in S$, there exists a maximum subgroup of $S$ containing $e$: the subset of $eSe$, the maximum subsemigroup of $S$ containing $e$ as its identity, of units with respect to $e$. We let ${\bf H}(e)$ be this maximum subgroup of $S$ containing idempotent $e$. Any subgroup of $S$ is a subgroup of the maximum subgroup containing its idempotent identity. For distinct idempotents $e$ and $f$, ${\bf H}(e)\cap{\bf H}(f)=\emptyset$. Let $E$ be the set of idempotents in $S$, $\{{\bf H}(e)|e\in E\}$ is the set of all maximal subgroups of $S$. These are also the $\mathcal{H}$-classes of $S$ containing idempotents.
@@ 4 Right [Left] Zero Subsemigroups
Let $S$ be a semigroup with idempotents $e$ and $f$. $e\,\mathcal{R}\,f$ if and only if ${\bf RZ}(e)={\bf RZ}(f)$.
By Proposition 4.3, $e\,\mathcal{R}\,f$ is equivalent to ${\bf RZ}(e)\cap{\bf RZ}(f)\neq\emptyset$. Similarly, for idempotents $e$ and $f$, $e\,\mathcal{L}\,f$ if and only if ${\bf LZ}(e)={\bf LZ}(f)$, and this is equivalent to ${\bf LZ}(e)\cap{\bf LZ}(f)\neq\emptyset$ by the dual of Proposition 4.3.
@@
If $e\,\mathcal{R}\,f$, then $eS^{1}=fS^{1}$. Since $e$ and $f$ are idempotents, $eS^{1}=eS$ and $fS^{1}=fS$. Thus $ex=f$ for some $x\in S$, and $e=fy$ for some $y\in S$. Then $eex=ef$, or $ex=ef$ because $e$ is an idempotent, and $f=ef$. Similarly, $fe=e$. By Proposition 4.3, ${\bf RZ}(e)={\bf RZ}(f)$.
If ${\bf RZ}(e)={\bf RZ}(f)$, then by Proposition 4.3, $ef=f$ and $fe=e$. Thus $efS=fS$ and $feS=eS$, so $fS\subset eS$ and $eS\subset fS$. Therefore $eS=fS$, which implies $eS^{1}=fS^{1}$ because $e$ and $f$ are idempotents, and $e\,\mathcal{R}\,f$. ∎
@@ 6 Rectangular Bands
In a band $S$ with element $a$, $V(a)=D_{a}$ and $a\,\mathscr{D}\,b$ if and only if ${\bf RZ}(a)\cap{\bf LZ}(b)\neq\emptyset$.
@@
Let $a,b\in S$ such that $c\in{\bf RZ}(a)\cap{\bf LZ}(b)\neq\emptyset$. Since $c$ is an idempotent, $c\in{\bf RZ}(c)\cap{\bf LZ}(c)$. Thus $c\in{\bf RZ}(a)\cap{\bf RZ}(c)$ and $c\in{\bf LZ}(c)\cap{\bf LZ}(b)$. Since these are nonempty, by Proposition 4.3 and Lemma 4.4 and their duals, $a\,\mathscr{R}\,c$ and $c\,\mathscr{L}\,b$, which implies $a\,\mathscr{D}\,b$. ∎
@@ 7 Commutative Bands
Let $S$ be a band. For any $a,b\in S$, if $\mathcal{LI}(a)=\mathcal{LI}(b)$ and $\mathcal{RI}(a)=\mathcal{RI}(b)$, then $a=b$.
@@
Since $S$ is a band, $a,b\in\mathcal{LI}(a)=\mathcal{LI}(b),\mathcal{RI}(a)=\mathcal{RI}(b)$. Therefore $ab=b$ because $a\in\mathcal{LI}(b)$, and $ab=a$ because $b\in\mathcal{RI}(a)$. Thus $a=b$. ∎
@@ 9 Commutative-rectangular Bands
  • if $T\neq\emptyset$, then $\mathcal{LZ}(c)=\displaystyle\cap_{x\in T}\mathcal{LZ}(x)$ and $\mathcal{RZ}(c)=\displaystyle\cap_{x\in T}\mathcal{RZ}(x)$, and
@@
Since $S$ is a commutative-rectangular band, either $a,b\in S$ are generalized inverses of each other or they commute. If $a$ and $b$ are generalized inverses of each other, then ${\bf RZ}(a)\cap{\bf LZ}(b)=T\neq\emptyset$ by Lemma 6.8. In this case, $aba=a$ and $bab=b$, and since $a$ and $b$ are idempotents, $aab=ab$ and $abb=b$. Therefore $ab\in T=\mathcal{RZ}(a)\cap\mathcal{LI}(a)\cap\mathcal{LZ}(b)\cap\mathcal{RI}(b)$. For any $x\in T$, $ax=x$, $xa=a$, $xb=x$, and $bx=b$, so that $abx=ab$ and $xab=ab$, meaning $ab\in\mathcal{LZ}(x),\mathcal{RZ}(x)$. Therefore $\mathcal{LZ}(ab)=\displaystyle\cap_{x\in T}\mathcal{LZ}(x)$ and $\mathcal{RZ}(ab)=\displaystyle\cap_{x\in T}\mathcal{RZ}(x)$. By the dual of Lemma 7.4, $ab$ must be the unique element of $S$ satisfying these one-sided zero sets.

### 17 Apr 2026 v4

@@ 3 Identity and Zero Relations and Sets
Let $S$ be a semigroup. For every idempotent $e\in S$, there exists a maximum subgroup of $S$ containing $e$: the subset of $eSe$, the maximum subsemigroup of $S$ containing $e$ as its identity, of units with respect to $e$. We let ${\bf H}(e)$ be this maximum subgroup of $S$ containing idempotent $e$. Any subgroup of $S$ is a subgroup of the maximum subgroup containing its idempotent identity. For distinct idempotents $e$ and $f$, ${\bf H}(e)\cap{\bf H}(f)=\emptyset$. Let $E$ be the set of idempotents in $S$, $\{{\bf H}(e)|e\in E\}$ is the set of all maximal subgroups of $S$. These are also the $\mathscr{H}$-classes of $S$ containing idempotents.
@@ 4 Right [Left] Zero Subsemigroups
Let $S$ be a semigroup with idempotents $e$ and $f$. $e\,\mathscr{R}\,f$ if and only if ${\bf RZ}(e)={\bf RZ}(f)$.
By Proposition 4.3, $e\,\mathscr{R}\,f$ is equivalent to ${\bf RZ}(e)\cap{\bf RZ}(f)\neq\emptyset$. Similarly, for idempotents $e$ and $f$, $e\,\mathscr{L}\,f$ if and only if ${\bf LZ}(e)={\bf LZ}(f)$, and this is equivalent to ${\bf LZ}(e)\cap{\bf LZ}(f)\neq\emptyset$ by the dual of Proposition 4.3.
@@
If $e\,\mathscr{R}\,f$, then $eS^{1}=fS^{1}$. Since $e$ and $f$ are idempotents, $eS^{1}=eS$ and $fS^{1}=fS$. Thus $ex=f$ for some $x\in S$, and $e=fy$ for some $y\in S$. Then $eex=ef$, or $ex=ef$ because $e$ is an idempotent, and $f=ef$. Similarly, $fe=e$. By Proposition 4.3, ${\bf RZ}(e)={\bf RZ}(f)$.
If ${\bf RZ}(e)={\bf RZ}(f)$, then by Proposition 4.3, $ef=f$ and $fe=e$. Thus $efS=fS$ and $feS=eS$, so $fS\subset eS$ and $eS\subset fS$. Therefore $eS=fS$, which implies $eS^{1}=fS^{1}$ because $e$ and $f$ are idempotents, and $e\,\mathscr{R}\,f$. ∎
Corollary 4.5.
Let $S$ be a band with element $e$. $R_{e}={\bf RZ}(e)$ and $L_{e}={\bf LZ}(e)$.
@@ 6 Rectangular Bands
In a band $S$ with elements $a$ and $b$, $V(a)=D_{a}$ and $a\,\mathscr{D}\,b$ if and only if ${\bf RZ}(a)\cap{\bf LZ}(b)\neq\emptyset$. When $a\,\mathscr{D}\,b$, ${\bf RZ}(a)\cap{\bf LZ}(b)=\{ab\}$.
@@
Let $a,b\in S$ such that $c\in{\bf RZ}(a)\cap{\bf LZ}(b)\neq\emptyset$. Since $c$ is an idempotent, $c\in{\bf RZ}(c)\cap{\bf LZ}(c)$. Thus $c\in{\bf RZ}(a)\cap{\bf RZ}(c)$ and $c\in{\bf LZ}(c)\cap{\bf LZ}(b)$. Since these are nonempty, by Proposition 4.3 and Lemma 4.4 and their duals, $a\,\mathscr{R}\,c$ and $c\,\mathscr{L}\,b$, which implies $a\,\mathscr{D}\,b$.
When $\mathscr{D}\,b, aa$ and $b$ are generalized inverses, so that $aba=a$ and $bab=b$, which implies $ab\in\mathcal{LI}(a)\cap\mathcal{RI}(b)$. Since $a$ and $b$ are idempotents, $aab=ab$ and $abb=ab$, which implies $ab\in\mathcal{RZ}\cap\mathcal{LZ}(b)$. Therefore $ab\in\mathcal{RZ}(a)\cap\mathcal{LI}(a)\cap\mathcal{LZ}(b)\cap\mathcal{RI}(b)={\bf RZ}(a)\cap{\bf LZ}(b)$. By Corollary 4.5, ${\bf RZ}(a)\cap{\bf LZ}(b)=R_{a}\cap L_{b}$. Since $S$ is a band, for any $e\in S$, $H_{e}={\bf H}(e)=\{e\}$, so $R_{a}\cap L_{b}=H_{ab}=\{ab\}$. ∎
@@ 7 Commutative Bands
Let $S$ be a groupoid. For any idempotents $a,b\in S$, if $\mathcal{LI}(a)=\mathcal{LI}(b)$ and $\mathcal{RI}(a)=\mathcal{RI}(b)$, then $a=b$.
@@
Since $a$ and $b$ are idempotents, $a,b\in\mathcal{LI}(a)=\mathcal{LI}(b),\mathcal{RI}(a)=\mathcal{RI}(b)$. Therefore $ab=b$ because $a\in\mathcal{LI}(b)$, and $ab=a$ because $b\in\mathcal{RI}(a)$. Thus $a=b$. ∎
@@ 9 Commutative-rectangular Bands
  • if $T\neq\emptyset$, then $\{c\}=T$, and
@@
Since $S$ is a commutative-rectangular band, either $a,b\in S$ are generalized inverses of each other or they commute. If $a$ and $b$ are generalized inverses of each other, then by Lemma 6.8, $\{ab\}={\bf RZ}(a)\cap{\bf LZ}(b)=T$.